Prepare 0.25M NaOH in 350ml (MW NaOH=40g); How about 0.1N NaOH?

Please show work
Answer #1

To make 0.25 M NaOH, you need to find out the required mass of NaOH.

Molarity = (given mass / molecular mass) × (1 / volume in litres)

0.25 = (given mass / 40) × (1000 / 350)

Given mass = 0.25 × 40 × 350 / 1000 = 3500 / 1000 = 3.5 g

So, you require 3.5 g NaOH to make 350 ml, 0.25 M of it.

For 0.1 N NaOH,

0.1 N = 0.1 M NaOH, it is because,

normality = n factor × molarity

n factor for NaOH is 1.

So, you need to prepare 0.1 M NaOH.

0.1 = (given mass / 40) × (1000 / 350)

Given mass = 0.1 × 40 × 350 / 1000 = 1.4 g

So, you require 1.4 g of NaOH to make 0.25 M, 350 ml of it.

Please rate.